OpenStudy (anonymous):

Find the requested composition of functions. Given f(x)=|16-x| and g(x)=3x+5, find (f . g)(x)

hero (hero):

(f o g)(x) = f(g(x)) f(g(x)) = |16 - (3x+5)| = |16 - 3x - 5|=|16 - 5 - 3x| = |11 - 3x|

OpenStudy (anonymous):

Thank-you! :D